# An image-computable model of speeded decision-making

### v1

@@ Methods
Target/flanker subspace alignment. To calculate the target/flanker subspace alignment metric, we first defined target and flanker subspaces using the SVM classifiers that we trained for the target/flanker decoding analyses. Specifically, for each of the four classifiers, which were trained to classify stimuli as a given target or flanker direction vs. all other target or flanker directions using the model features from a given layer, we extracted the vector orthogonal to the decision hyperplane. In agreement with prior work [4, 40], we refer to these vectors as decoding vectors for a given target/flanker direction. Let $\mathbf{x}_{k,\textrm{targ}}^{T}$ and $\mathbf{x}_{k,\textrm{flnk}}^{T}$ denote the target decoding row vector and flanker decoding row vector for the $k$th direction, respectively. We define the matrices formed with these four decoding vectors filling the rows as the target subspace matrix $\mathbf{X}_{\textrm{targ}}$ and the flanker subspace matrix $\mathbf{X}_{\textrm{flnk}}$. These matrices have dimensions $4\times K_{l}$, where $K_{l}$ is the number of feature dimensions for layer $l$. Each matrix therefore spans a subspace of the full $K_{l}$-dimensional feature space. Our goal is to determine whether the target and flanker subspaces are orthogonal.
@@
To do so, we employ principal angles between subspaces [88, 31], which generalizes the more intuitive notion of angles between lines or planes to arbitrary dimensions. To calculate the principal angles, we require orthonormal bases for the target and flanker subspaces, which we determine using a reduced singular value decomposition (SVD):
@@
Participation ratio. We measured the dimensionality of target representations for a given layer with the participation ratio ($\textrm{PR}_{l}$) [21], defined as:
@@
where $K_{l}$ is the number of features in layer $l$ and $\lambda_{1}\geq\ldots\geq\lambda_{i}\geq\ldots\geq\lambda_{K_{l}}$ are the eigenvalues of the target-centered feature covariance matrix for layer $l$. The target-centered features were obtained by subtracting the centroid of the feature matrix for each target direction from the corresponding trials in the feature matrix [56].
The participation ratio is a continuous measure of dimensionality ranging from 1 to $K_{l}$. The minimum ($\textrm{PR}_{l}=1$) is obtained when all of the feature variance is concentrated in a single dimension, such that $\lambda_{i}=0$ for $i\geq 2$. The maximum ($\textrm{PR}_{l}=K_{l}$) is obtained when the variance is evenly spread across the $K_{l}$ feature dimensions, such that all $K_{l}$ eigenvalues are equal.
@@
The mutual information for a given stimulus feature was normalized by the entropy of the feature distribution to facilitate comparisons between the features [47]. The entropy is given by:
@@ Acknowledgements
Fig. S4: RT delta plots and conditional accuracy functions. A) RT delta plots for participants and models trained with the standard training paradigm (all RT data; $n$ = 75 models/participants). B) Conditional accuracy functions for participants and models trained with the standard training paradigm. C) RT delta plots for participants and models trained separately on data in each RT quantile, then combined ($n$=10 combined models/participants). D) Conditional accuracy functions for participants and models trained separately on data in each RT quantile, then combined. For all panels, error bars correspond to bootstrap 95% confidence intervals.

### v2

@@ Methods
Target/flanker subspace alignment. To calculate the target/flanker subspace alignment metric, we first defined target and flanker subspaces using the SVM classifiers that we trained for the target/flanker decoding analyses. Specifically, for each of the four classifiers, which were trained to classify stimuli as a given target or flanker direction vs. all other target or flanker directions using the CNN activations from a given layer, we extracted the vector orthogonal to the decision hyperplane. In agreement with prior work [bernardi2020abstraction, libby2021orthogonal], we refer to these vectors as decoding vectors for a given target/flanker direction. Let $\mathbf{x}_{k,\textrm{targ}}^{T}$ and $\mathbf{x}_{k,\textrm{flnk}}^{T}$ denote the target decoding row vector and flanker decoding row vector for the $k$th direction, respectively. We define the matrices formed with these four decoding vectors filling the rows as the target subspace matrix $\mathbf{X}_{\textrm{targ}}$ and the flanker subspace matrix $\mathbf{X}_{\textrm{flnk}}$. These matrices have dimensions $4\times K_{l}$, where $K_{l}$ is the number of active units in layer $l$. Each matrix therefore spans a subspace of the full $K_{l}$-dimensional space. Our goal is to determine whether the target and flanker subspaces are orthogonal.
@@
To do so, we employ principal angles between subspaces [zhu2013principal_angles, jordan1875principal_angles], which generalizes the more intuitive notion of angles between lines or planes to arbitrary dimensions. To calculate the principal angles, we require orthonormal bases for the target and flanker subspaces, which we determine using a reduced singular value decomposition (SVD):
@@
Participation ratio. We measured the dimensionality of target representations for a given layer with the participation ratio ($\textrm{PR}_{l}$) [gao2017_participation_ratio], defined as:
@@
where $K_{l}$ is the number of active units in layer $l$ and $\lambda_{1}\geq\ldots\geq\lambda_{i}\geq\ldots\geq\lambda_{K_{l}}$ are the eigenvalues of the target-centered activation covariance matrix for layer $l$. The target-centered activations were obtained by subtracting the centroid of the activation matrix for each target direction from the corresponding trials in the activation matrix [rangamani2023neural_collapse].
The participation ratio is a continuous measure of dimensionality ranging from 1 to $K_{l}$. The minimum ($\textrm{PR}_{l}=1$) is obtained when all of the variance in activity is concentrated in a single dimension, such that $\lambda_{i}=0$ for $i\geq 2$. The maximum ($\textrm{PR}_{l}=K_{l}$) is obtained when the variance is evenly spread across the $K_{l}$ dimensions, such that all $K_{l}$ eigenvalues are equal.
@@
The mutual information for a given stimulus feature was normalized by the entropy of the feature distribution to facilitate comparisons between the features [muratore2022prune]. The entropy is given by:
@@ Acknowledgements
Fig. S4: RT delta plots and conditional accuracy functions. A) RT delta plots for participants and VAMs ($n$ = 75 models/participants). B) Conditional accuracy functions for participants and VAMs. For all panels, error bars correspond to bootstrap 95% confidence intervals.